OpenStudy (sar9):

Two angles of a triangle have the same measure and the third one is 3 degrees greater than the measure of each of the other two. Find the measure of the LARGEST angle in the triangle

OpenStudy (kittiwitti1):

\[\angle1=\angle2\ne\angle3;~\angle3=3^{o}+\angle1,2\]\[\angle1+\angle2+\angle3=180^{o}\]

OpenStudy (kittiwitti1):

let's put angles 1 and 2 as A and B, and angle 3 as C. Since A=B let's just use A.\[A+A+C\rightarrow C=3+A\rightarrow A+A+(3+A)=180\]

OpenStudy (kittiwitti1):

Or...\[3A+3=180\]So...\[3A=177\]\[A=\frac{177}{3}\]Since C is the greatest angle value, just add 3 to that.\[C=\frac{177}{3}+3=?\]